OpenStudy (anonymous):

21. graph of the function y = –1/3 × 4^x

OpenStudy (dumbcow):

First off , you need to need that horizontal asymptote for standard exponential functions is y=0 then just plug in a couple points, x=0 and x=1 are the best to use f(0) = -1/3 f(1) = -4/3 connect the points on a graph and make the line approach y=0 done:)